OpenStudy (anonymous):

A circle is drawn on a coordinate plane so that the endpoints of a diameter are located on points (-1,4) and (7,2). What is the area of the circle?

OpenStudy (anonymous):

1. Find the distance from (-1,4) to (7,2) using the distance formula 2. The radius of your circle is 1/2 of that distance 3. Area = Pi * radius^2

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ 4}})\quad &({\color{red}{ 7}}\quad ,&{\color{blue}{ 2}}) \end{array}\qquad diameter = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2} \\ \quad \\ \textit{area of a circle}=\pi r^2\qquad \qquad r=\frac{diameter}{2}\)